overview of the aiv test campaign n.
Download
Skip this Video
Loading SlideShow in 5 Seconds..
Overview of the AIV test campaign PowerPoint Presentation
Download Presentation
Overview of the AIV test campaign

play fullscreen
1 / 21

Overview of the AIV test campaign

157 Views Download Presentation
Download Presentation

Overview of the AIV test campaign

- - - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript

  1. Overview of the AIV test campaign Willem Luinge Consortium meeting, SVR #2, TRR December 2006, Groningen

  2. Outline • Achievements since SVR • Status of integration • Pre ILT • EGSE/ICC/test infra structure/test scripts • Test equipment • Miscellaneous • Outlook at the near future • Full ILT (radiometry, stability, spectral tests) in the last test weeks

  3. Integration steps

  4. WU rack

  5. Integration activities of HIFI units (1) Mechanical, electrical, command, optical, RF, thermal interfaces to HIFI parts and test equipment; performance check of back ends • WBS, Dec 2005. • Recent IF tuning algorithm works better • HRS integration (10 – 14 April; HRS-H last wk of July) • Performance re-test of back ends hindered by other integrations and Testcontrol problems. Most recent IF stability test are being evaluated.

  6. Integration activities of HIFI units (2) • ICU/OBS integration (ICU: June) • ICU: Boot procedure ok for ILT (OBSM not in place) • Test OBS 4.3 and various units is ongoing; almost all functions for ILT implemented; maintenance still needed • OBS FT: open • FCU FM integration (early July and early November) • Upgrade of power line filtering is implemented (ok with SFT) • Re-qual is open • FPU integration (early Aug) • Was used as detector for LOU (Aug) • Beam measurements complete (Sep) • Open: Chopper scan, fast chop mode, IF feedback, mixer magnet tuning algorithm, generation of diplexer tables, linearity test to be completed/evaluated, chopper response time, thermal settling times • Update and re-run of Integration and FT scripts

  7. Integration activities of HIFI units (3) • LCU FM integration (started 2nd week August) • Integration ICU - LCU FM mechanical, electrical integration and low level command interface tested including electrical interface LCU LSU dummy and LOA dummy • Procedure to upload LCU safety table upload and memory patch operational (various releases tables, MIB and CUS needed) (end of November!) • Passed LO SFT (dummy) • Testmodes for LO ss tested • FM upconverter integration (Oct.), ok. Performance untested. • LSU FM is delivered, TM properties were measured (voltage, IF out put vs. freq and LSU settings) • EIDP’s of FP and LO

  8. Integration activities of HIFI units (4) • LOU FM integration (started 2nd week August) • LOU installed mechanically, checked, cooled, checked again and aligned visually • Check mixer pump level by LOU (controlled by LCU IMD2 and SCOE) for most bands  presentation Nick Whyborn • Bands 3a, 7a did not pass integration test (during last cons meeting) • Dismount LO bands & inspection, analysis of failed bands, etc • Nov/Dec: remount other LOA’s and test (IMD2, SCOE), integrate with LCU FM, LSU FM  LO presentation • Open: RF integration (pump level, noise temperature) with LSU FM

  9. Set up End of November

  10. Pre ILT • Beam measurements of all bands before & after vibration: complete except for analysis and evaluation (dedicated talk) • Linearity tests band 1 are to be evaluated (beam splitter adapted) • Testmode Mixer magnet tuning result differs from the result of a slow scan of magnet current in band 4 (bands 2, 3, 5: open) • IF Stability test: results not fully understood, investigation is not finished • Integration and functional test: are being reviewed taking into account the lessons learned. Then a re-run is needed before delivery • Status gas cell pre ILT tests

  11. EGSE/ICC/test infra structure/test scripts • Testcontrol • Time jitter problems due to undefined time steps in SCOS release and CDMS simulator (1553 bus protocol) • Work around by inserting delays between the commands (flight efficiency cannot be tested) • Bottlenecks from HCSS side seem to have disappeared (TM ingest and other) • New HCSS version makes data almost on line available • A new release has to be made after every change in configuration (MIB, CUS, config file value) • MIB, CUS, config file are in line with current integration phase • New scripts are debugged and tested in parallel set up as far as configuration allows • HK verify scripts ready for test on main line • DB access/IA/know how to use IA is not widespread

  12. Open items Test equipment • Stable operation of FPU cryostat so far, extra safety cycle required? Supply of (liquid) helium? • Programming TEI’s re-imager, PDU // set up and H/C source: to be modified/untested • Matching optics re-imager fit checked (pupil-to-field) alignment to be improved

  13. Miscellaneous issues • Safety issues are in place • No major change in manpower situation, analysis groups installed • Documentation • ILT procedure are split; need editing • Activity sheets added • Modification of the Integration and FT will have an impact on the IST procedures

  14. Outlook • See planning • Integration not completed (set back e.g. LO problems) • Next: complete LO integration • [Pre ILT gas cell measurements] • ILT (=Radiometry) for current bands can start provided • Mixer magnet tuning is in place • Diplexer tables are available • Chopper scan confirms the current assumptions • Other test (settling time etc, chopper response time, fast chopmode) to fill open slots • January: spectral tests and gas cell tests • February: refurbishment etc • March: Band 3, 7 tests • Stability tests and operational procedures mainly overnight

  15. De volgende sheets zijn spares

  16. Set up August 2006

  17. Test engineers HIFI ILT

  18. Man power support FM ILT campaign • Mechanical integration: RvdS, ME, LM, WH • Operation, installation of LO sources: NW, WJ, PD, MJ • Alignment issues: BK, MJ, RH, ME • Test equipment handling (vacuum & cryogens): WH, FW • Electrical integration: HS, AN, HG • EGSE, MIB: LD (back up NB, AdJ) • Operators: NB, RH, (MB), back up: AdJ • ICC (software infrastructure, database, IA, etc): AdJ, KE, PZ and others • CUS scripts: DT • Analysis & reporting: Testengineers, system engineer, calibration group, others • General: Contact persons for various subsystems

  19. Documents • Reference documents, VCM • AIV plan • Detailed AIV plan (FM ILT): under configuration control after this review • HIFI ILT Integration and Functional Test Procedure (calls for the scripts): under configuration control after this review • “HIFI ILT Performance Test Procedure” • FM ILT CUS scripts, including documents (e.g. a description: CUS scripts for HIFI FM tests), configuration files • MIB(’s) • IA use cases document and IA tools • Various integration, configuration procedures, manuals • A list of websites was added to the TRR data package • Issue for TRR: content of procedures for ILT and approach